# Object-Attribute-Relation Representation based Video Semantic Communication

### v1

@@ IV-C Performance Evaluation of OAR-assisted JSCC
Figure 7: The image perceptual quality (LPIPS) at different CBRs and SNRs. (a)-(b) illustrate the LPIPS versus SNR at different CBRs, and (c)-(d) illustrate the LPIPS versus CBR at different SNRs.
Figure 8: Performance comparison of OAR-based video transmission to H.265 paired with LDPC and DVST. (a)-(b) are performances at CBR of 1/300, and (c)-(d) at SNR of 0 dB. The H.265-based transmission under some parameters suffer from decode failure and is not plotted.
@@ IV-C1 OAR-assisted Image JSCC for Reference Frame Transmission
Performance of OAR-assisted JSCC is compared with BPG-based transmission and base JSCC for reference frame transmission. Fig. 7(a)-(b) depict the image LPIPS variation with SNR at different CBRs, while Fig. 7(c)-(d) illustrate the variation with the CBR at different SNRs. OAR-assisted JSCC generally outperforms base JSCC, particularly at high CBR (Fig. 7(b)) and low SNR (Fig. 7(c)). Specifically, compared to most BPG+LDPC+QAM schemes, OAR-assisted JSCC avoids cliff effects. It also surpasses BPG+1/3LDPC+4QAM and BPG+1/2LDPC+BPSK methods, which do not suffer from cliff effects in the 0$\sim$20 dB interval, particularly at high SNR (Fig. 7(b)). Furthermore, Fig. 7(c) demonstrates that OAR-assisted JSCC still maintains effective transmission at low SNR, while most BPG-based transmissions encounter truncation, failing to achieve effective transmission. Simultaneously, Fig. 7(d) highlights the superior performance of OAR-assisted JSCC at higher SNRs.
@@ IV-C2 OAR-assisted Video Transmission through JSCC
The overall performance of the proposed video transmission system, integrating OAR-based video coding and OAR-assisted JSCC, is assessed. Figure 8 compares the performance of the proposed OAR-based video transmission system with traditional H.265-based mechanisms and the deep JSCC method DVST[9]. Fig. 8(a) and (b) demonstrate the variations in perceptual quality (LPIPS) and downstream task performance (mAP) at a CBR of 1/300. The proposed OAR-based method demonstrates superior perceptual quality and task performance, particularly at low SNRs. While H.265-based video transmission exhibits better quality at higher SNRs, it suffers from the cliff effect. Notably, under a CBR constraint of 1/300, the proposed method consistently outperforms DVST across all SNRs, achieving an average LPIPS loss reduction of 0.037 and mAP enhancement of 0.16.
Fig. 8(c) and (d) illustrate the variation in performances with CBR at a fixed SNR of 0 dB. It is evident that the proposed method achieves optimal perceptual quality and downstream task performance in the low CBR range (below 0.006). This is attributed to the capability to effectively reconstruct objects by leveraging explicit semantics of OAR and prior knowledge of the generative model. The performance improvement of our proposed method is comparatively gradual with increasing CBR compared to other methods. This is because other methods achieve video quality enhancement by transmitting residuals. Hence, our method is better suitable for environments with severely limited bandwidth and significant noise interference.
@@ V Analysis, Discussions and Ablation Studies
V-A Appearance Enhancement with Semantic Priors
Experiments conducted on OAR-based video coding at low bit-rates reveal that the proposed video generation model effectively enhances reconstruction of objects using semantic priors. Illustrated in Fig. 9, this effect is demonstrated by showing the 5th, 10th, and 15th frames of the reconstructed video alongside their corresponding local details. It is evident that the reconstruction model progressively incorporates prior information from the training set into the reconstructed frames, resulting in enhanced object details.
Figure 9: The video reconstruction results with reference frames of poor quality. The reference frame is taken as CRF 48.
TABLE I: Variation of mAP achieved by warmup operation
@@ V-A Appearance Enhancement with Semantic Priors
Accordingly, we propose a warm-up strategy to further validate the effectiveness of the prior knowledge. Specifically, we use the compressed first frame as the reference frame and use the corresponding OAR for image generation. The above process is repeated for several iterations, with each generated image serving as the reference frame for the next iteration, which is called the warm-up stage. Subsequent frames are then generated after completing the warm-up stage. Table I presents the variations in mAP with warm-up of 10 iterations on the val1 and val2 datasets. It is observed that warm-up at low bit-rates enhances the video quality to a certain extent. However, when the reference frames exhibit good quality (with CRF values of 42 and above), excessive inference by the model results in the loss of original details, leading to degradation of video quality.
V-B Ablation Studies for OAR Attributes and Relations
To assess the influence of each element in OAR, ablation experiments are conducted across three dimensions: object categories, orientations, and relations. In each experiment, the same model is utilized for video generation, with different alterations to the OAR sequences: (1) w/o category: setting all categories to “car”; (2) w/o angle: fixing all angles to 0; (3) w/o relation: removing all occlusion relations. The performance variation in mAP is shown in Table II. As indicated from the table, object category and angle produce significant impact on reconstruction performance, whereas relation has minimal impact. This suggests that occlusion relations may be implicitly embedded in spatial locations due to the independence of objects, rendering them redundant in the OAR. However, the inclusion of relations in OARs ensures the integrity of the graph structure, facilitating the adaptation and expansion of OAR representations to more complex scenes.
Figure 10: Visualization of ablation experiments. (a) Real image; (b) Reference frame; (c) Original reconstructed frame; (d) Reconstructed frame of ablation setting: w/o category; (e) Reconstructed frame of ablation setting: w/o angle. The bottom halves of (d) and (e) are the same as those in (c).
TABLE II: Average variations in mAP metrics of videos obtained by different ablation methods
@@ V-B Ablation Studies for OAR Attributes and Relations
Fig. 10 demonstrates further video reconstruction results from the ablation experiments (1) and (2). As shown in Fig. 10 (a) to (c), the van absent in the reference frame is generated leveraging category information from the OAR, and is correctly recognized as a “van” with high confidence. This indicates the adaptability of OAR-based video transmission to downstream tasks. While in Fig. 10(d), mislabeling the category as “car” results in car-like appearances, such as a black body and a smaller front end. However, the bus in the upper-left corner, primarily influenced by appearance information from the reference frame, remains unaffected by OAR mislabeling. This reflects the capability of OAR to explicitly govern video semantics, suggesting its potential application in diverse domains like video editing and privacy protection. Fig. 10(e) demonstrates the rectification of the model for angle mislabeling. This suggests the angle’s impact on object appearance is not as pronounced as category, manifesting instead through a composite effect with background or other features. On the other hand, this also indicates that the object orientation is also concurrently influenced by motion.
@@
From the above analysis, it is evident that the attributes of object categories have explicit influence on appearance. To delve deeper into the interplay between categories, the impact of category mislabeling on various objects is examined. Experiments show that the average precision of “van” decreases by 0.316 and 0.340 on val1 and val2, respectively, while “bus” and “others” have relatively minor variations.
@@
Objects categorized as “bus” and “others” exhibit less susceptibility to category mislabeling due to their distinct sizes and movement patterns compared to cars. Conversely, vans are most affected by mislabeling, with nearly half of originally correctly recognized objects (with mean mAPs of 0.643 and 0.484 for val1 and val2, respectively) being misidentified. Further experiments reveal that after being mislabeled as car, the proportion of vans misidentified as cars increases from 35% to 67%. This indicates that objects with similar appearances are more susceptible to category labeling. This further exemplifies the explicit role of OAR in semantics and inspires awareness of distinctiveness among different attributes during labeling design.
V-C Performance Variants for OAR-assisted JSCC
Fig. 11 compares the LPIPS versus SNR performances of OAR-assisted JSCC with the base JSCC. At low CBRs (e.g., CBR=1/24), OAR-assisted JSCC does not outperform base JSCC, particularly at high SNRs. However, as CBR increases, OAR-assisted JSCC surpasses base JSCC, presenting a significant advantage. This is attributed to the increased number of output channels in the network, allowing more OAR prior information to be reflected in transmitted symbols and facilitating efficient decoding at the receiver. In fact, given that downstream tasks prioritize perceptual and semantic considerations, it is more valuable to optimize the loss function from perceptual and semantic perspectives [44, 45]. However, since this paper primarily focuses on video transmission with OAR as auxiliary information, refining the JSCC model specifically for reference frames falls outside the scope of this study.
@@ V-C Performance Variants for OAR-assisted JSCC
Figure 11: The PSNR versus SNR curves of the foreground region obtained by different JSCC models for the three CBR values.
@@ VI Conclusion
In this paper, an OAR-assisted video coding and transmission method was proposed to address the challenge of efficient and reliable video transmission for downstream tasks. First, an OAR-based video representation and coding system was proposed to realize a low bit-rate representation of videos by object-attribute-relation (OAR). Second, a generative video reconstruction method based on reference frames and OAR sequences was implemented to realize perceptual-quality and downstream-task oriented reconstruction. Finally, OAR-assisted JSCC was proposed and combined with OAR-based video coding to realize video transmission in noisy channels. Experiments on a traffic surveillance dataset demonstrated the effectiveness of our approach. First, the proposed OAR-based video coding outperformed H.265 in terms of perceptual quality and downstream object detection performance and saved up to 61% bit-rate. Second, the implemented OAR-based video transmission system could be adaptable to low bandwidth and SNR conditions, effectively ensuring perceptual quality and downstream task performance. At a CBR of 1/300, our method achieved an average LPIPS loss reduction of 0.037 and object detection performance (mAP) improvement of 0.16 compared to DVST. Additionally, the scalability of the proposed OAR representation suggests its potential applicability in complex scenarios such as border and sea defense, as well as applications requiring explicit semantic control like privacy-preserving video transmission.

### v2

@@ IV-C1 OAR-assisted Image JSCC for Reference Frame Transmission
The performance of OAR-assisted JSCC is compared with BPG-based transmission and base JSCC for reference frame transmission on the UA-DETRAC dataset. Fig. 9(a)-(b) depict the image LPIPS variation with the SNR at different CBRs, and Fig. 9(c)-(d) illustrate the variation with the CBR at different SNRs. OAR-assisted JSCC generally outperforms base JSCC in general. Specifically, compared with most BPG+LDPC+QAM schemes, OAR-assisted JSCC avoids cliff effects. It also outperforms the BPG+1/3LDPC+4QAM and BPG+1/2LDPC+BPSK methods, which do not suffer from cliff effects in the 0$\sim$20 dB interval, particularly at high SNRs (Fig. 9(b)). Furthermore, Fig. 9(c) demonstrates that OAR-assisted JSCC still maintains effective transmission at low SNRs, whereas most BPG-based transmissions encounter truncation, which fails to achieve effective transmission.
@@ IV-C2 OAR-assisted Video Transmission through JSCC
The overall performance of the proposed video transmission system, which integrates OAR-based video generation and OAR-modulated JSCC, is assessed. Fig. 10 compares the performance of the proposed OAR-based video transmission system with traditional H.265-based mechanisms and the deep JSCC method DVST[9] on the UA-DETRAC dataset. Fig. 10(a) and (b) demonstrate the variations in perceptual quality (LPIPS) and downstream task performance (mAP) at a CBR of 1/300. The proposed OAR-based method demonstrates superior perceptual quality and task performance, particularly at low SNRs. While H.265-based video transmission exhibits better quality at higher SNRs, it suffers from the cliff effect. Notably, under a CBR constraint of 1/300, the proposed method consistently outperforms DVST across all SNRs, achieving an average LPIPS loss reduction of 0.037 and an mAP enhancement of 0.16.
Fig. 10(c) and (d) illustrate the variation in performances with respect to the CBR at a fixed SNR of 0 dB. The proposed method achieves optimal perceptual quality and downstream task performance in the low CBR range (below 0.006). This is attributed to the ability to effectively reconstruct objects by leveraging the explicit semantics of OAR and prior knowledge of the generative model. Compared with those of the other methods, the performance improvement of our proposed method is relatively gradual with increasing CBR. This is because other methods achieve video quality enhancement by transmitting residuals. Hence, our method is better suited for environments with severely limited bandwidth and significant noise interference.
Video transmission performances on the CATER and SoccerNet datasets are also evaluated from the same perspectives and presented in Fig. 11 and Fig. 12, respectively. The proposed OAR-based video transmission system maintains high perceptual quality and task performance even at low CBRs and low SNRs. This indicates the important role of explicit semantic information contained in OAR. Even though a low CBR and low SNR introduce a large distortion for the transmission of key frames, the video generator leverages the semantic prior to achieve faithful video synthesis based on OAR. This demonstrates the important advantage of the proposed framework for the transmission of semantically distinct scenarios involving objects with simple appearances.
@@ V Analysis, Discussions and Ablation Studies
V-A Ablation study on the deformation branch in OAR-based video generation
To avoid redundancy, only the results of the experiments conducted on the UA-DETRAC dataset are presented below. The CATER and SoccerNet datasets yield similar conclusions. In the proposed video reconstruction framework shown in Fig. 3, frame prediction is achieved via a two-branch fusion method: the synthesis branch (producing $\tilde{\bf I}_{t}$) and the deformation branch (producing $\check{\bf I}_{t}$). The final reconstructed image is obtained through a weighted summation of the outputs from the two branches. The deformation branch performs optical flow estimation based on the OAR layout of sequential frames and is used to synthesize images consistent with the motion characteristics. We quantitatively evaluate the impact of the deformation branch in terms of pixel-level distortion, perceptual distortion, and task performance. The results are presented in Table I.
TABLE I: Results of the ablation study on the OAR-based video generation networks
PSNR LPIPS mAP
two branch ($\hat{\bf I}_{t}$) 19.73 0.389 0.565
two branch ($\tilde{\bf I}_{t}$) 19.33 0.401 0.540
single branch ($\hat{\bf I}_{t}$) 18.62 0.438 0.498
The comparison between the synthesized image $\tilde{\bf I}_{t}$ and the final output image $\hat{\bf I}_{t}$ in the two-branch structure demonstrates that the introduction of the deformation branch significantly enhances the reconstruction performance. The reconstructed video after two-branch fusion shows noticeable performance gains compared with using only the synthesis branch. Furthermore, when the deformation branch is removed and the entire video reconstruction relies solely on the synthesis branch, substantial performance degradation is observed across all the metrics. This highlights the effectiveness of the deformation branch in improving video quality, particularly for long-term video synthesis.
V-B Appearance Enhancement with Semantic Priors
Experiments conducted on OAR-based video coding at low bit-rates reveal that the proposed video generation model effectively enhances the reconstruction of objects via semantic priors. As illustrated in Fig. 13, this effect is demonstrated by showing the 5th, 10th, and 15th frames of the reconstructed video alongside their corresponding local details. The reconstruction model progressively incorporates prior information from the training set into the reconstructed frames, resulting in enhanced object details.
Accordingly, we propose a warm-up strategy to further validate the effectiveness of prior knowledge. Specifically, we take the compressed first frame as the reference frame and use the corresponding OAR for image generation. The above process is repeated for several iterations, with each generated image serving as the reference frame for the next iteration, which is called the warm-up stage. Subsequent frames are then generated after completing the warm-up stage. Table II presents the variations in mAP with warm-up of 10 iterations on the val1 and val2 datasets. Warm-up at low bit-rates enhances the video quality to a certain extent. However, when the reference frames exhibit good quality (with CRF values of 42 and below), excessive inference by the model results in the loss of original details, leading to degradation of video quality.
Figure 13: Video reconstruction results with reference frames of poor quality. The reference frame is taken as CRF 48.
TABLE II: Variation in the mAP achieved by the warmup operation
@@ V-B Appearance Enhancement with Semantic Priors
V-C Ablation Studies for OAR Attributes and Relations
To assess the influence of each element in the OAR, ablation experiments are conducted across three dimensions: object categories, orientations, and relations. In each experiment, the same model is utilized for video generation, with different alterations to the OAR sequences: (1) w/o category: all categories are set to “car”; (2) w/o angle: all angles to fixed to 0; (3) w/o relation: all occlusion relations are removed. The performance variation in mAP is shown in Table III. As indicated in the table, the object category and angle significantly affect the reconstruction performance, whereas the relation has a minimal impact. This suggests that occlusion relations may be implicitly embedded in spatial locations owing to the independence of objects, rendering them redundant in the OAR. However, the inclusion of relations in OARs ensures the integrity of the graph structure, facilitating the adaptation and expansion of OAR representations to more complex scenes. Moreover, relations in OAR are able to convey high-level semantic information within a very small amount of data, which is conducive to the development of video transmission frameworks for both reconstruction and downstream tasks. Specifically, the receiver can perform downstream tasks either through direct reasoning based on the OAR or by using the OAR as auxiliary information for scene understanding. The relations in the OAR complement the behavioral logic and interactions between key entities in the scene at the semantic level, information that may not be effectively preserved in the reconstructed video under low bit-rate or low SNR conditions. An important direction for our future research is to explore how to further utilize the relations in OAR to achieve efficient video transmission and downstream task execution across a broader range of scenarios.
Figure 14: Visualization of ablation experiments. (a) Real image; (b) Reference frame; (c) Original reconstructed frame; (d) Reconstructed frame with the ablation setting: w/o category; (e) Reconstructed frame with the ablation setting: w/o angle. The bottom halves of (d) and (e) are the same as those in (c).
TABLE III: Average variations in mAP metrics of videos obtained via different ablation methods
@@ V-C Ablation Studies for OAR Attributes and Relations
Fig. 14 shows further video reconstruction results from ablation experiments (1) and (2). As shown in Fig. 14 (a) to (c), the van absent in the reference frame is generated leveraging category information from the OAR, and is correctly recognized as a “van” with high confidence. This indicates the adaptability of OAR-based video transmission to downstream tasks. In Fig. 14(d), mislabeling the category as “car” results in car-like appearances, such as a black body and a smaller front end. However, the bus in the upper-left corner, which is influenced primarily by appearance information from the reference frame, remains unaffected by OAR mislabeling. This reflects the ability of OAR to explicitly govern video semantics, suggesting its potential application in diverse domains, such as video editing and privacy protection. Fig. 14(e) shows the rectification of the model for angle mislabeling. This suggests that angle’s impact on object appearance is not as pronounced as that of the category, manifesting instead through a composite effect with background or other features. On the other hand, this also indicates that the object orientation is also concurrently influenced by motion.
@@
From the above analysis, it is evident that the attributes of object categories have explicit influence on appearance. To delve deeper into the interplay between categories, the impact of category mislabeling on various objects is examined. Experiments show that the average precision of “van” decreases by 0.316 and 0.340 on val1 and val2, respectively, whereas “bus” and “others” have relatively minor variations.
@@
Compared with cars, objects categorized as “bus” and “others” are less susceptible to category mislabeling because of their distinct sizes and movement patterns. Conversely, vans are most affected by mislabeling, with nearly half of the originally correctly recognized objects (with mean mAPs of 0.643 and 0.484 for val1 and val2, respectively) being misidentified. Further experiments reveal that after being mislabeled as a car, the proportion of vans misidentified as cars increases from 35% to 67%. This finding indicates that objects with similar appearances are more susceptible to category labeling. This further exemplifies the explicit role of OAR in semantics and inspires awareness of the distinctiveness among different attributes during labeling design.
@@
Figure 15: PSNR versus SNR curves of the foreground region obtained by different JSCC models for the three CBR values.
V-D Performance Variants for OAR-assisted JSCC
Fig. 15 compares the LPIPS versus SNR performances of OAR-assisted JSCC with the base JSCC. The experimental results indicate that the JSCC performance under different CBRs and SNRs is significantly improved by introducing OAR modulation, and an LPIPS loss reduction of up to 0.023 can be achieved. This finding indicates that the explicit semantic information contained in OAR plays an effective role in assisting the transmission and reconstruction of key frames. From the results of CBR=1/24 and 1/16, it is evident that the modulation effect of the OAR is dependent on a better channel environment. When transmitting with a lower CBR, a higher channel SNR is more helpful for the efficient expression of semantic information in OAR. Since downstream tasks prioritize perceptual and semantic considerations, it is more valuable to optimize the loss function from perceptual and semantic perspectives [57, 58]. However, since this paper focuses primarily on video transmission with OAR as auxiliary information, refining the JSCC model specifically for reference frames falls outside the scope of this study.
V-E Inference time analysis
The inference time of several key modules of the proposed framework was measured experimentally at two different image resolutions, and the results are presented in Table IV. The tested modules include JSCC encoding and decoding of key frames, OAR extraction, OAR embedding and graph computation, and image generation. The experiments were conducted on an A800 GPU. It should be noted that the current system focuses on the proposal of the OAR-based video transmission system and its rate-semantic-distortion performance, while computational complexity is not the primary focus.
TABLE IV: Inference time for different modules
module
inference time/ms
($512\times 512$)
inference time/ms
($256\times 256$)
JSCC Encoding 31.20 14.72
JSCC Decoding 6.92 7.44
OAR Extraction 9.8 1.4
OAR Embedding
and Graph Computing
18.80 9.22
Image Generation 22.35 9.18
Total 89.07 41.96
However, various optimization techniques can enhance inference speed without significantly compromising video quality. In terms of network design, smaller models can be achieved by reducing the number of convolutional kernels, convolutional layers, and embedding dimensions, thereby accelerating model inference. Additionally, techniques such as knowledge distillation and pruning can be employed to create lightweight models. Model efficiency and inference speed can also be improved through parameter quantization and low-precision inference. Furthermore, pipeline design and parallel processing of multiple frames offer potential acceleration methods. However, these optimizations fall outside the scope of this paper and will be addressed in our future research.
@@ VI Conclusion
In this paper, an OAR-assisted video coding and transmission method was proposed to address the challenge of efficient and reliable video transmission for downstream tasks. First, an OAR-based video representation and coding system was proposed to realize a low bit-rate representation of videos via object-attribute-relation (OAR). Second, a generative video reconstruction method based on reference frames and OAR sequences was implemented to realize perceptual-quality and downstream-task oriented reconstruction. Finally, OAR-assisted JSCC was proposed and combined with OAR-based video coding to realize video transmission in noisy channels. Experiments on a traffic surveillance dataset demonstrated the effectiveness of our approach. First, the proposed OAR-based video coding outperformed H.265 in terms of perceptual quality and downstream object detection performance and saved up to 61% bit-rate. Second, the implemented OAR-based video transmission system could be adaptable to low bandwidth and SNR conditions, effectively ensuring perceptual quality and downstream task performance. At a CBR of 1/300, our method achieved an average LPIPS loss reduction of 0.054 and an improvement in the object detection performance (mAP) of 0.15 compared with DVST. Additionally, the scalability of the proposed OAR representation suggests its potential applicability in complex scenarios such as border and sea defenses, as well as in applications requiring explicit semantic control such as privacy-preserving video transmission.